# Skeleton-in-Context: Unified Skeleton Sequence Modeling with In-Context Learning

### v1

@@ 5 Experiments
Implementation Details. We implement the proposed Skeleton-in-Context model with the number of layers $N$ = 5, number of attention heads $H$ = 8, and hidden feature dimension $C$ = 256. For each prompt input/target and query input/target, the sequence length is $F$ = 16. All tasks are unified into a one-off, end-to-end training process, without any task-specific designs. The training takes about 6 hours for 120 epochs on 4 NVIDIA GeForce RTX 4090 GPUs.
@@ 5.2 Compared with SoTA Methods
Quantitative Results. We report experimental results on four tasks in Tab. 1. We term Skeleton-in-Context implemented with static/dynamic-TUP as Static/Dynamic-SiC, respectively. For other models, we reproduce them and report the results. As can be seen from Tab. 1, our proposed model outperforms all multi-task models, achieving state-of-the-art results on the in-context multi-task benchmark. Notably, the results of task-specific models on a single task reflect the difficulty of each task. However, we also exhibit impressive performance in comparison with task-specific models, which demonstrates our proposed model is an expert in handling skeleton sequence-based multi-tasking.
@@ 5.4 Ablation Study
Prompt Engineering. Since previous work has proven that different prompts in context learning affect the performance of the model, we conduct ablation experiments on the choice of prompts. As Tab. 4 shows, we select the prompt by the feature similarity and MPJPE minimum between the input skeleton sequence and the prompt respectively. Experimental results show that the random selection method achieves the best results. Unlike 2D images and 3D point clouds, skeleton-based task data is highly similar, resulting in different prompts playing similar roles in the same task. While the random selection method expands the model’s receptive field to the data and achieves better results.
@@
Architecture. We conduct ablation experiments on the architecture of the model. As shown in Tab. 5, after exploring different hidden feature dimensions and transformer depths, we find that the model performs best when the feature dimension is $256$ and the depth is $5$.
@@
5.5 In-Context Prior Pose
We conduct an in-depth exploration of Dyanamic-TUP. We extract the trained Dyanamic-TUP weight parameters, then randomly initialize a $(F,J,3)$ tensor and use the trained encoder to increase its feature dimension to C-dimension. Finally, an optimizer is used to optimize the tensor and minimize its feature similarity with the trained Dyanamic-TUP. The top of Fig. 6 represents the visualization of Static-TUP, while the bottom is the Dyanamic-TUP. The concrete skeleton of Dyanamic-TUP is similar to the mean pose Static-TUP of the training set, which demonstrates that the Dyanamic-TUP learned by our model is actually the prior knowledge whose receptive field is all ground truth in the training set. This prior knowledge helps our model learn more smoothly on multiple tasks, and adding learnable modules helps our model generalize to other datasets or new tasks (Sec. 5.3).
@@ 6 Conclusion
We propose the Skeleton-in-Context, designed to process multiple skeleton-base tasks simultaneously after just one time of training. Specifically, We build a skeleton-based in-context benchmark covering a wide range of tasks. Moreover, We propose skeleton prompts composed of TGP and TUP, which solve the overfitting problem of skeleton sequence data trained under the training framework commonly applied in previous 2D and 3D in-context models. Besides, we demonstrate that our model can generalize to different datasets and new tasks, such as motion completion. We hope it paves the way for further exploration of in-context learning in the skeleton-based sequences.
Supplementary Material
{strip}
Table 6: Importance of our proposed task-guided prompt (TGP) and task-unified prompt (TUP) employed in Skeleton-in-Context (SiC). TGP and TUP are important designs as they address the challenge of multi-task training. Please refer to Sec. B for more discussion.
Methods MP. (AMASS) PE. (H3.6M) JC. (3DPW) FPE. (H3.6M)
[…]
Task-Specific Model: one architecture for one task
SiC (w/o TGP&TUP) 57.7 46.2 41.6 36.7 47.7 46.0 56.5 45.4 49.7 59.8 54.7 62.5 72.6 67.6
SiC (w/o TGP&TUP) 62.9 49.8 44.3 36.3 46.5 47.9 59.1 47.4 51.3 61.6 56.5 76.4 86.8 81.6
Static-SiC 11.0 19.0 22.7 32.6 37.8 24.6 51.6 41.8 29.8 44.7 37.2 59.8 67.2 63.5
Dynamic-SiC 11.0 18.9 22.6 32.2 37.4 24.4 51.8 42.5 29.5 44.0 36.8 59.2 66.5 62.9
Overview. The supplementary material includes sections as follows:
  • Section A presents more details on our experiments and a more detailed description of the datasets.
  • Section B verifies the effectiveness of our Skeleton-in-Context on multi-task synergistic training.
  • Section C provides more detailed results both quantitatively and qualitatively.
@@ A.1 Implementation Details
We implement the proposed Skeleton-in-Context model with the number of layers $N$ = 5, number of attention heads $H$ = 8, and hidden feature dimension $C$ = 256. For each prompt input/target and query input/target, the sequence length is $F$ = 16. We implement Skeleton-in-Context with PyTorch. In the default setting, during both training and evaluation, for each query pair, we randomly select a prompt pair from the training set of the same task as the query pair. We use an AdamW optimizer with a linearly decaying learning rate, starting at 0.0002 and decreasing by 1% after every epoch. All tasks are unified into a one-off, end-to-end training process, without any task-specific designs. The training takes about 6 hours for 120 epochs on 4 NVIDIA GeForce RTX 4090 GPUs.
@@ A.2 Datasets and Metrics
H3.6M (Human3.6M) is used for the pose estimation and future pose estimation tasks. It is a large-scale dataset, which contains 3.6 million video frames of actions involving 15 types and 7 actors. Following previous works [76], we use subjects 1, 5, 6, 7, and 8 for training, and subjects 9, 11 for testing, and preprocessing the poses. After preprocessing, each pose has 17 joints. We use the Stacked Hourglass (SH) networks [43] to extract the 2D skeletons from videos as input. For pose estimation, we expect the model to estimate the corresponding 3D skeletons. For future pose estimation, we expect the model to predict and estimate at the same time the future 3D skeletons in a time range of 300ms, given the history 300ms of 2D skeletons. As we use the same dataset on two tasks, it makes multi-tasking harder as the model needs further context to correctly perceive and then accomplish the task. In Skeleton-in-Context we use prompt to guide the model to learn in context.
AMASS is used for the motion prediction task. It integrates most of the existing marker-based Mocap datasets, which are parameterized with a unified representation. We follow common practices in human motion prediction [40, 18] to use AMASS-BMLrub for testing and the rest of the datasets for training. For motion prediction, we expect the model to predict the future 3D motion sequence given the history 3D motion sequence. The time ranges of the future and history motion sequences are both 400ms, which is 10 frames under the frame rate of 25 fps. As the model requires the sequence to be of 16 frames, we pad the last poses 6 times.
3DPW (3D Pose in the Wild) is used for the joint completion task. It has more than 51k frames with 3D annotations for challenging indoor and outdoor activities. For joint completion, we construct 2 settings, where we respectively randomly mask 40% and 60% of all the joints. We expect the model to reconstruct the missing joints.
@@ B.1 Negative Transfer in Multi-Task
A straight way to train a generalist multi-task model is to collect the data from multiple tasks, format them into the same shape, and directly train the model on them. However, as explained in [62], a phenomenon named negative transfer in multi-dataset training will occur and lead to poor results in the above training method. As different tasks have their own unique goals, they may confuse the model during training without guidance from context, leading to performance drop in testing. To demonstrate the effect of negative transfer, as shown in Tab. 6, we train and test the backbone of our SiC separately on four tasks, whose results are in Tab. 6 with gray background. Note that the backbone here does not include TUP and TGP. When we train the four tasks together, the performance of the model decreases due to the data variability between the individual tasks, which can be attributed to the negative transfer phenomenon [62]. The experiments show that, without context guidance from prompts (TUP and TGP), the multi-task training is limited by negative transfer and not able to achieve satisfactory performance.
@@ B.2 Multi-Task Synergistic Training
Our Skeleton-in-Context enables multiple tasks to be trained in a unified way, we term it Multi-Task Synergistic Training. With our proposed Task-Guided Prompt (TGP), SiC can perceive context (the corresponding task-specific information) from TGP, and accomplish the task accurately. Meanwhile, in order to deal with data from different datasets and tasks, which have different patterns and may affect the performance if not processed properly, we introduce the Task-Unified Prompt (TUP) as prior knowledge of the query target. As a task-agnostic module, the TGP is able to encode the unified human motion representations of various tasks. As Tab. 6 shows, with the proposed TGP and TUP, our SiC is able to achieve impressive results on four tasks simultaneously.

### v2

@@ 5 Experiments
Implementation Details. We implement the proposed Skeleton-in-Context model with the number of layers $N$ = 5, number of attention heads $H$ = 8, and hidden feature dimension $C$ = 256. For each prompt input/target and query input/target, the sequence length is $F$ = 16. All tasks are unified into a one-off, end-to-end training process without any task-specific designs. The training takes about 6 hours for 120 epochs on 4 NVIDIA GeForce RTX 4090 GPUs.
@@ 5.2 Compared with SoTA Methods
Quantitative Results. We report experimental results on four tasks in Tab. 1. We term Skeleton-in-Context implemented with static/dynamic-TUP as Static/Dynamic-SiC, respectively. For other models, we reproduce them and report the results. As can be seen from Tab. 1, our proposed model outperforms all multi-task models, achieving state-of-the-art results on the in-context multi-task benchmark. Notably, the results of task-specific models on a single task reflect the difficulty of each task. However, we also exhibit impressive performance compared to task-specific models, demonstrating that our proposed model is an expert in handling skeleton sequence-based multi-tasking.
@@ 5.4 Ablation Study
Prompt Engineering. Since previous work has proven that different prompts in context learning affect the performance of the model, we conduct ablation experiments on the choice of prompts. As Tab. 4 shows, we select the prompt by the feature similarity and MPJPE minimum between the input skeleton sequence and the prompt, respectively. Experimental results show that the random selection method achieves the best results. Unlike 2D images and 3D point clouds, skeleton-based task data is highly similar, resulting in different prompts playing similar roles in the same task. While the random selection method expands the model’s receptive field to the data and achieves better results.
@@
Architecture. We conduct ablation experiments on the architecture of the model. As shown in Tab. 5, after exploring different hidden feature dimensions and transformer depths, we find that the model performs best when the feature dimension is $256$, and the depth is $5$.
@@
5.5 In-Context Initial Pose
We conduct an in-depth exploration of Dynamic-TUP. We extract the trained Dynamic-TUP weights, then randomly initialize a $(F,J,3)$ tensor and use the trained encoder to increase its feature dimension to C-dimension. Finally, an optimizer is used to optimize the tensor and minimize its feature similarity with the trained Dynamic-TUP. The top of Fig. 6 represents the visualization of Static-TUP, while the bottom is the Dynamic-TUP. The concrete skeleton of dynamic-TUP is similar to the mean pose Static-TUP of the training set, which demonstrates that the Dynamic-TUP learned by our model is actually the prior knowledge whose receptive field is all ground truth in the training set. This prior knowledge helps our model learn more smoothly on multiple tasks, and adding learnable modules helps our model generalize to other datasets or new tasks (Sec. 5.3).
@@ 6 Conclusion
In this work, we propose the Skeleton-in-Context, designed to process multiple skeleton-base tasks simultaneously after just one training time. Specifically, we build a novel skeleton-based in-context benchmark covering various tasks. In particular, we propose skeleton prompts composed of TGP and TUP, which solve the overfitting problem of skeleton sequence data trained under the training framework commonly applied in previous 2D and 3D in-context models. Besides, we demonstrate that our model can generalize to different datasets and new tasks, such as motion completion. We hope our research builds the first step in the exploration of in-context learning for skeleton-based sequences, which paves the way for further research in this area.
Acknowledgement This work is supported by National Natural Science Foundation of China (No. 62203476), Natural Science Foundation of Shenzhen (No. JCYJ20230807120801002). This project is also supported by the interdisciplinary doctoral grants (iDoc 2021-360) from the Personalized Health and Related Technologies (PHRT) of the ETH domain, Switzerland.
@@ A.1 Implementation Details
We implement the proposed Skeleton-in-Context model with the number of layers $N$ = 5, number of attention heads $H$ = 8, and hidden feature dimension $C$ = 256. For each prompt input/target and query input/target, the sequence length is $F$ = 16. We implement Skeleton-in-Context with PyTorch. In the default setting, during both training and evaluation, for each query pair, we randomly select a prompt pair from the training set of the same task as the query pair. We use an AdamW optimizer with a linearly decaying learning rate, starting at 0.0002 and decreasing by 1% after every epoch. All tasks are unified into a one-off, end-to-end training process without any task-specific designs. The training takes about 6 hours for 120 epochs on 4 NVIDIA GeForce RTX 4090 GPUs.
@@ A.2 Datasets and Metrics
H3.6M (Human3.6M) is used for the pose estimation and future pose estimation tasks. It is a large-scale dataset that contains 3.6 million video frames of actions involving 15 types and 7 actors. Following previous works [79], we use subjects 1, 5, 6, 7, and 8 for training and subjects 9 and 11 for testing and preprocessing the poses. After preprocessing, each pose has 17 joints. We use the Stacked Hourglass (SH) networks [42] to extract the 2D skeletons from videos as input. For pose estimation, we expect the model to estimate the corresponding 3D skeletons. For future pose estimation, we expect the model to predict and estimate at the same time the future 3D skeletons in a time range of 300ms, given the history 300ms of 2D skeletons. As we use the same dataset on two tasks, multi-tasking is harder as the model needs further context to correctly perceive and then accomplish the task. In Skeleton-in-Context we use prompts to guide the model to learn in context.
AMASS is used for the motion prediction task. It integrates most of the existing marker-based Mocap datasets, which are parameterized with a unified representation. We follow common practices in human motion prediction [39, 19] to use AMASS-BMLrub for testing and the rest of the datasets for training. For motion prediction, we expect the model to predict the future 3D motion sequence, given the 3D history motion sequences. The time ranges of the future and history motion sequences are both 400ms, which is 10 frames under the frame rate of 25 fps. As the model requires the sequence to be of 16 frames, we pad the last poses 6 times.
3DPW (3D Pose in the Wild) is used for the joint completion task. It has more than 51k frames with 3D annotations for challenging indoor and outdoor activities. For joint completion, we construct 2 settings, where we randomly mask 40% and 60% of all the joints. We expect the model to reconstruct the missing joints.
@@ B.1 Negative Transfer in Multi-Task
A straight way to train a generalist multi-task model is to collect the data from multiple tasks, format them into the same shape, and directly train the model on them. However, as explained in [65], a phenomenon named negative transfer in multi-dataset training will occur and lead to poor results in the above training method. As different tasks have their own unique goals, they may confuse the model during training without guidance from context, leading to a performance drop in testing. To demonstrate the effect of negative transfer, as shown in Tab. 6, we train and test the backbone of our SiC separately on four tasks, whose results are in Tab. 6 with gray background. Note that the backbone here does not include TUP and TGP. When we train the four tasks together, the performance of the model decreases due to the data variability between the individual tasks, which can be attributed to the negative transfer phenomenon [65]. The experiments show that, without context guidance from prompts (TUP and TGP), the multi-task training is limited by negative transfer and not able to achieve satisfactory performance.
@@ B.2 Multi-Task Synergistic Training
Our Skeleton-in-Context enables multiple tasks to be trained in a unified way, which we term Multi-Task Synergistic Training. With our proposed Task-Guided Prompt (TGP), SiC can perceive context (the corresponding task-specific information) from TGP, and accomplish the task accurately. Meanwhile, in order to deal with data from different datasets and tasks, which have different patterns and may affect the performance if not processed properly, we introduce the Task-Unified Prompt (TUP) as prior knowledge of the query target. As a task-agnostic module, the TGP is able to encode the unified human motion representations of various tasks. As Tab. 6 shows, with the proposed TGP and TUP, our SiC is able to achieve impressive results on four tasks simultaneously.